OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y < -x + 6? (-3, 0) (-1, 0) (4, 3) (3, -2)

OpenStudy (anonymous):

Plug the values in. 0 < 3 + 6 = 9 YES 0 < 1 + 6 = 7 YES 3 < -4+6 = 2 NO -2 < -3 + 6 = 3 YES

OpenStudy (anonymous):

so the third one :)